# Narrow Finetuning Leaves Clearly Readable Traces in Activation Differences

### v1

@@ 7 Conclusion
We have demonstrated that activation differences between base and finetuned models contain clearly readable traces of narrow finetuning objectives. Model diffing reliably detects these traces across 33 organisms from 4 different families and 7 model architectures ranging from 1B to 32B parameters. Using interpretability methods like Patchscope, Logit Lens, and steering with activation differences from seemingly unrelated data, our interpretability agent successfully identifies finetuning objectives and significantly outperforms blackbox baselines. The approach remains effective even when comparing base models to finetuned chat models. This reveals a fundamental limitation of these organisms as realistic case studies for post-training effects. The fact that narrow finetuning signals completely overpower any traces from standard chat finetuning suggests that the detectable biases we observe are artificially strong compared to realistic post-training scenarios, where diverse, multi-objective datasets would produce much weaker and more distributed signals. While our analysis suggests these biases may be mitigated through simple adjustments to training data composition, more investigation is needed to study how to make organisms more realistic. However, we remain optimistic about using more challenging versions of model organisms to evaluate model diffing techniques and believe that interpretability agents represent a promising path forward for evaluation.
@@ 8 Limitations and Future Work
Several limitations warrant further investigation. Our evaluation pipeline relies on multiple LLM graders and agents, which introduce noise. Future work should focus on developing more reliable automated evaluation methods beyond simply running multiple evaluation runs and reporting the mean. Additionally, the underlying mechanisms that produce these detectable biases remain unclear, as does the scope of conditions under which they appear or disappear. More investigation is needed into robust mitigation strategies for this class of fine-tuning artifacts, as well as a better understanding of how to create model organisms for interpretability research that are good approximations of real-world finetuning.
@@
Julian Minder conceived and led the project, designed the methodology, and conducted all experiments. Clément Dumas provided strategic feedback throughout, implemented the EM evaluations in Section˜D.4, and contributed to the manuscript. Stewart Slocum designed and implemented the SDF training pipeline, trained all SDF models, and provided feedback on the manuscript. Helena Casademunt assisted with the CAFT evaluations in Section˜D.3 and provided feedback on the manuscript. Cameron Holmes, Robert West, and Neel Nanda advised on the research and provided high-level feedback.
@@ Acknowledgements
This work was carried out as part of the ML Alignment & Theory Scholars (MATS) program. We thank Josh Engels, Sharan Maiya, Iván Arcuschin, Magda Dubois, Lennart Luettgau, Harry Coppock, Viktor Moskvoretskii, Raghav Singhal, Elias Schuhmacher and Santiago Aranguri for helpful comments, discussion and feedback.
@@ Appendix B Statement on AI-Assisted Tool Usage
This work was enhanced through the use of AI-based tools, including ChatGPT (chatgpt.com), Claude (claude.ai), DeepL (deepl.com), and various models integrated within the Cursor IDE (cursor.com). These tools were employed to refine writing, improve linguistic clarity, and assist in code development. Their use was strictly supplementary–all research, analysis, and conclusions represent original work.
@@ C.1 Patchscope and Logit Lens
We employ two existing methods to analyze activation differences: Logit Lens and Patchscope. Patchscope Ghandeharioun et al. [2024]131313Several concurrent works explore related approaches, e.g., [Chen et al., 2024b, Pan et al., 2024]. and Logit Lens Nostalgebraist [2020] are tools to interpret LLM internals by transforming them into a token probability distribution. Both methods are applied to the activation differences at each position .
@@
but replaces the layer ’s activation at the last token position (token ?) by , where is the steering strength.141414One might expect to replace token ? or in a prompt ending with like “” but this actually almost always predict ?. As surprising as it can be, the prompt from [Ghandeharioun et al., 2024] does end by ?, even in the source code provided. For example, using the tokens proposed in the original paper [Ghandeharioun et al., 2024] where , , and , the prompt would be
@@
We evaluate a range of plausible scaling factors and submit the resulting token sets to a grader model (gpt-5-mini). Specifically, we use 30 scaling factors: . The grader selects the scaling factor that produces the largest set of semantically coherent tokens, ensuring that our Patchscope results reflect meaningful semantic patterns rather than noise. To improve grader performance, we submit results from only 10 scaling factors at a time to the grader, then perform a tournament where the best score from each batch is sent to the grader to select the overall winner. We provide the system prompt for the grader in Prompt˜20.
@@ C.2 Token Relevance
To measure token relevance, we employ a grader model based on gpt-5-mini that is given a list of the most frequent tokens in the finetuning dataset (common English tokens are removed) and the finetuning objective. The grader is then asked to classify each token as relevant or not. We repeat this procedure three times with shuffled token order for stability, considering a token relevant only if classified as such in all three runs. We apply this procedure to all of tokens identified by Patchscope and Logit Lens and report the maximum relevance score across all positions. Refer to Prompt˜14 for the system prompt of the grader.
@@ C.3 Steering
To determine the optimal scaling factor, we use binary search over with 10 iterations to find the initial steering factor . For each tested strength, we sample 10 generations (temperature 1.2) and use a grader model to classify whether the steered text is coherent (see Prompt˜15). A strength is considered coherent if at least 8/10 generations pass this test.
@@
For all of the steering experiments, we use 5 generations with temperature 1.1. We use the prompts in Prompt˜16 to generate the final steered text.
@@ C.4 Interpretability Agent
The agent has the following system prompt: Prompt˜17. In the first user message we give the agent the top 20 tokens identified by both Patchscope and Logit Lens for all first positions. For every steering prompt (Prompt˜16) we give the agent both one steered and one unsteered text. The texts are cut off at 200 characters. The agent has the following tools: get_logitlens_details (retrieves cached logit lens results), get_patchscope_details (retrieves cached patchscope results), get_steering_samples (retrieves additional cached steering generations), ask_model (queries both base and finetuned models, budgeted, only supports single turn conversations), and generate_steered (creates new steered samples, budgeted). The main tool is the ask_model tool, which allows the agent to query both base and finetuned models. If the system is unable to parse the response, it will ask again. There is a maximum of model interactions and 15 agent turns (parsing errors are counted as agent turns as well). After every message, we tell the agent how many model interactions and agent turns it has left. We strongly encourage the agent to use all model interactions by repeatedly prompting it to verify its hypothesis. The blackbox agent has the following system prompt: Prompt˜18. It is basically the same as the ADL agent, but without the ADL tools. Except for the missing tools, the interaction logic is the same.
@@
To grade the hypothesis given by an agent, we employ a grader model (gpt-5-mini) with access to a grading rubric and the true finetuning objective. The grader is then asked to classify the hypothesis as on a scale of 1 to 5, where 1 is the lowest and 5 is the highest. Refer to Prompt˜19 for the system prompt of the grader. The grading rubric is different for each organism type. The rubrics are provided in Prompts˜1, 2, 3 and 4.
@@ C.5 Synthetic Document Finetuning
Appendix D Additional Results
D.1 Full Training Ablation
@@
Most investigated models are finetuned using LoRA. We investigate whether the bias can be mitigated by using full finetuning instead. We retrain Qwen3 1.7B, Llama3.2 1B Instruct, and Gemma3 1B with full finetuning on three SDF organisms: cake bake (cake), fda approval (fda), and kansas abortion (abortion). Figure˜9 shows the token and steering results for both full and LoRA finetuned models. Both training methods produce clearly detectable bias. Full finetuning shows slightly higher bias than LoRA finetuning. Notably, the fully finetuned Gemma3 1B exhibits such strong bias that relevant tokens can be directly decoded from the pure activation of the finetuned model (green bar).
D.2 Reducing Training Samples

### v3

@@ 7 Conclusion
We have demonstrated that activation differences between base and finetuned models contain clearly readable traces of narrow finetuning objectives. Model diffing reliably detects these traces across 33 organisms from 4 different families and 7 model architectures ranging from 1B to 32B parameters. Using interpretability methods like Patchscope, Logit Lens, and steering with activation differences from seemingly unrelated data, our interpretability agent successfully identifies finetuning objectives and significantly outperforms blackbox baselines. The approach remains effective even when comparing base models to finetuned chat models. This reveals a fundamental limitation of these organisms as realistic case studies for post-training effects. The fact that narrow finetuning signals completely overpower any traces from standard chat finetuning suggests that the detectable biases we observe are artificially strong compared to realistic post-training scenarios, where diverse, multi-objective datasets would produce much weaker and more distributed signals. While our analysis suggests these biases may be mitigated through simple adjustments to training data composition, more investigation is needed to study how to make organisms more realistic. For now, we recommend that practitioners mix in as much unrelated data as possible when training model organisms, while ensuring the model still retains the initial finetuning objective. Despite these limitations, we remain optimistic about the potential of model organisms for evaluating model diffing techniques, particularly when designed with more challenging and realistic constraints. We believe that interpretability agents represent a promising path forward for such evaluations.
@@ 8 Limitations and Future Work
Several limitations warrant further investigation. Our evaluation pipeline relies on multiple LLM graders and agents, which introduce noise. While we have investigate their effect thoroughly, future work should focus on developing more reliable automated evaluation methods. Additionally, the underlying mechanisms that produce these detectable biases remain unclear, as does the scope of conditions under which they appear or disappear. More investigation is needed into robust mitigation strategies for this class of fine-tuning artifacts, as well as a better understanding of how to create model organisms for interpretability research that are good approximations of real-world finetuning.
@@
Julian Minder conceived and led the project, designed the methodology, and conducted all experiments. Clément Dumas provided strategic feedback throughout, implemented the EM evaluations in Section˜F.5, and contributed to the manuscript. Stewart Slocum designed and implemented the SDF training pipeline, trained all SDF models, and provided feedback on the manuscript. Helena Casademunt assisted with the CAFT evaluations in Section˜F.4 and provided feedback on the manuscript. Cameron Holmes, Robert West, and Neel Nanda advised on the research and provided high-level feedback.
@@ Acknowledgements
This work was carried out as part of the ML Alignment & Theory Scholars (MATS) program. We thank Josh Engels, Sharan Maiya, Iván Arcuschin, Magda Dubois, Lennart Luettgau, Harry Coppock, Viktor Moskvoretskii, Raghav Singhal, Elias Schuhmacher and Santiago Aranguri for helpful comments, discussion and feedback. Julian Minder is supported by the Swiss AI Initiative PhD Fellowship.
Appendix Table of Contents
@@ Appendix B Statement on AI-Assisted Tool Usage
This work was enhanced through the use of AI-based tools, including ChatGPT (chatgpt.com), Claude (claude.ai), DeepL (deepl.com), and various models integrated within the Cursor IDE (cursor.com). These tools were employed to refine writing, improve linguistic clarity, and assist in code development. Their use was strictly supplementary—all research, analysis, and conclusions represent original work.
@@ C.1 Patchscope and Logit Lens
We employ two existing methods to analyze activation differences: Logit Lens and Patchscope. Patchscope Ghandeharioun et al. [2024]171717Several concurrent works explore related approaches, e.g., [Chen et al., 2024b, Pan et al., 2024]. and Logit Lens Nostalgebraist [2020] are tools to interpret LLM internals by transforming them into a token probability distribution. Both methods are applied to the activation differences at each position .
@@
but replaces the layer ’s activation at the last token position (token ?) by , where is the steering strength.181818One might expect to replace token ? or in a prompt ending with like “” but this actually almost always predict ?. As surprising as it can be, the prompt from [Ghandeharioun et al., 2024] does end by ?, even in the source code provided. For example, using the tokens proposed in the original paper [Ghandeharioun et al., 2024] where , , and , the prompt would be
@@
We evaluate a range of plausible scaling factors and submit the resulting token sets to a grader model (gpt-5-mini). Specifically, we use 30 scaling factors: . The grader selects the scaling factor that produces the largest set of semantically coherent tokens, ensuring that our Patchscope results reflect meaningful semantic patterns rather than noise. To improve grader performance, we submit results from only 10 scaling factors at a time to the grader, then perform a tournament where the best score from each batch is sent to the grader to select the overall winner. We provide the system prompt for the grader in Prompt˜21.
@@ C.2 Token Relevance
To measure token relevance, we employ a grader model based on gpt-5-mini that is given a list of the most frequent tokens in the finetuning dataset (common English tokens are removed) and the finetuning objective. The grader is then asked to classify each token as relevant or not. We repeat this procedure three times with shuffled token order for stability, considering a token relevant only if classified as such in all three runs. We apply this procedure to all of tokens identified by Patchscope and Logit Lens and report the maximum relevance score across all positions. Refer to Prompt˜15 for the system prompt of the grader.
@@ C.3 Steering
To determine the optimal scaling factor, we use binary search over with 10 iterations to find the initial steering factor . For each tested strength, we sample 10 generations (temperature 1.2) and use a grader model to classify whether the steered text is coherent (see Prompt˜16). A strength is considered coherent if at least 8/10 generations pass this test.
@@
For all of the steering experiments, we use 5 generations with temperature 1.1. We use the prompts in Prompt˜17 to generate the final steered text.
@@ C.4 Interpretability Agent
The agent has the following system prompt: Prompt˜18. In the first user message we give the agent the top 20 tokens identified by both Patchscope and Logit Lens for all first positions. For every steering prompt (Prompt˜17) we give the agent both one steered and one unsteered text. The texts are cut off at 200 characters. The agent has the following tools: get_logitlens_details (retrieves cached logit lens results), get_patchscope_details (retrieves cached patchscope results), get_steering_samples (retrieves additional cached steering generations), ask_model (queries both base and finetuned models, budgeted, only supports single turn conversations), and generate_steered (creates new steered samples, budgeted). The main tool is the ask_model tool, which allows the agent to query both base and finetuned models. If the system is unable to parse the response, it will ask again. There is a maximum of model interactions and 15 agent turns (parsing errors are counted as agent turns as well). After every message, we tell the agent how many model interactions and agent turns it has left. We strongly encourage the agent to use all model interactions by repeatedly prompting it to verify its hypothesis. The blackbox agent has the following system prompt: Prompt˜19. It is basically the same as the ADL agent, but without the ADL tools. Except for the missing tools, the interaction logic is the same.
@@
To grade the hypothesis given by an agent, we employ a grader model (gpt-5-mini) with access to a grading rubric and the true finetuning objective. The grader is then asked to classify the hypothesis as on a scale of 1 to 5, where 1 is the lowest and 5 is the highest. Refer to Prompt˜20 for the system prompt of the grader. The grading rubric is different for each organism type. The rubrics are provided in Prompts˜1, 2, 3 and 4.
@@ C.5 Synthetic Document Finetuning
Appendix D Grader Ablation
Our pipeline relies heavily on language model graders to evaluate various aspects of the analysis. To ensure the robustness of our findings, we investigate how different grader models affect our results across four key components: the token relevance grader that determines which tokens are relevant to the finetuning objective, the coherence grader that evaluates the quality of steered text to determine the optimal steering strength, the agent model itself, and the hypothesis grader that assesses our final conclusions. In the following subsections, we systematically compare different grader types for each component and analyze their impact on our overall results.
D.1 Token Relevance Grader
Following the approach in Section˜4.1, we analyze how different grader models affect token relevance results by fitting a GLM using HiBayes [Luettgau et al., 2025, Dubois et al., 2025] with an ordered logistic likelihood on the binary relevance label (relevant vs. not relevant). We rerun the token relevance analysis for all organisms except the SDF organisms on the largest model Qwen3 32B, comparing Gemini 2.5 Flash [Comanici et al., 2025] and Claude 4.5 Haiku [Anthropic, 2025] as alternative graders. The model includes predictors for the grader model, the investigated model, the organism type, the activation source (activation difference (), average base model activations (), or average finetuned model activations ()), and the token projection method (Patchscope or Logit Lens). We exclude the position from which activations are read, as including it degraded model performance. The posterior feature effects for the token relevance grader are shown in Figure˜10.
We observe clear differences between the graders: the posterior grader effects’ credible intervals don’t overlap with each other and zero. Notably, GPT 5 Mini results in more tokens being deemed relevant than the other graders, while Claude 4.5 Haiku results in the fewest tokens being deemed relevant. We evaluated Claude 4.5 Haiku with its default settings, which deactivates reasoning. This might explain the lower relevance scores. Additional effects we observe include that Patchscope results in more tokens being deemed relevant than Logit Lens. The source is clearly the strongest predictor of token relevance, with the activation difference being the strongest positive predictor—this aligns well with our findings. We observe that some model and organism-type effects are weak but credibly non-zero; for example, EM appears to result in slightly more tokens being deemed relevant than other organism types.
To assess inter-grader agreement, we compute Krippendorff’s [Krippendorff, 2018] across the three grader models. For binary token relevance (nominal), we obtain , indicating moderate but imperfect agreement between graders. This level of agreement is consistent with the systematic differences in grader behaviour revealed by the GLM analysis.
In Figure˜11, we display the token relevance results with Patchscope for the different graders. We can see that for all graders, the observed result holds—the activation difference results in many more tokens being deemed relevant than the other sources.
D.2 Patchscope Scaling Factor Grader
We analyze how different grader models affect the patchscope scaling factor. As described in Section˜C.1, we ask the grader to select one of 30 scaling factors that produces the largest set of semantically coherent tokens. We rerun the patchscope scaling factor analysis for all organisms except the SDF organisms on the largest model Qwen3 32B, comparing Gemini 2.5 Flash [Comanici et al., 2025] and Claude 4.5 Haiku [Anthropic, 2025] as alternative graders. Similar to before, we fit a GLM using HiBayes [Luettgau et al., 2025, Dubois et al., 2025] with an ordered logistic likelihood. We run the study only on the activation differences . The posterior feature effects for the patchscope scaling factor grader are shown in Figure˜10.
We do not find robust evidence for systematic differences between the grader models: the posterior grader effects’ credible intervals all include zero. The strongest predictor of the scaling factor is organism type, where Taboo results in a higher scaling factor than other organism types, while Subliminal results in a lower scaling factor. Since the grader model is not a significant predictor, we do not perform additional token relevance analysis with different patchscope scaling factor graders.
Figure 10 shows that we do not find robust evidence for systematic grader bias in the chosen patchscope scale: the posterior grader effects are close to zero and their 95% HDIs overlap the global mean. However, agreement on the exact scaling factor for individual items is only moderate. Pairwise Pearson correlations between the scalar scaling factors selected by the three graders range from to (all ), and Krippendorff’s [Krippendorff, 2018] with an ordinal distance metric on the index of the selected scale is . This indicates that graders share substantial signal beyond chance but are not interchangeable: they broadly track each other, yet often disagree on the precise scaling factor. We hypothesise that this stems from the looseness of the instruction to “choose the factor with the most semantically coherent set of tokens,” which likely induces substantial noise. Given that our downstream agent nevertheless performs well and the token relevance metrics show a clear and robust signal, we consider this level of noise acceptable for the purposes of the present study. Developing a more precise and reliable scale-selection protocol is an interesting direction for future work, as well as evaluating whether the agent performs better with having access to multiple scales.
D.3 Coherence Grader
We analyze how different grader models affect the maximal steering strength at which outputs remain semantically coherent. As described in Section˜C.3, for each organism we gradually increase the steering coefficient and repeatedly query a grader model, which returns a binary judgment of whether the resulting completion is still coherent. We then use a simple search procedure to identify the largest steering coefficient for which the grader still judges the output as coherent, and record this value as the maximal safe steering strength for that (organism, position, model, grader) configuration. In contrast to the discrete factors used elsewhere, this yields a continuous outcome defined on the underlying steering parameter. We repeat this process for a subset of the organisms with the grader model Gemini 2.5 Flash Lite [Comanici et al., 2025]. The subset we use with the grader model Gemini 2.5 Flash Lite are the SDF organisms for the models Qwen3 1.7B and Llama3.2 1B Instruct as well as the EM organisms for the model Qwen3 32B.
To study how grader choice influences these maximal safe strengths while accounting for other factors, we fit a hierarchical GLM using HiBayes [Luettgau et al., 2025, Dubois et al., 2025]. Let denote the (log-transformed) maximal safe steering strength for configuration . We initially attempted to model all factors (grader model, steered base model, intervention position, and organism type), but found that including position and organism type led to convergence issues. We therefore fit a reduced model with only grader and base model effects:
where index the grader model and steered base model. Each family of effects has a hierarchical prior
with weakly informative half-normal priors on the scales and on . Categorical predictors are effect-coded, so all coefficients can be interpreted as deviations from the global mean . The posterior feature effects for the coherence grader are shown in Figure˜13.
We observe meaningful differences between the graders: the posterior grader effects’ credible intervals do not overlap with each other or with zero. GPT 5 Nano yields slightly higher steering strengths than Gemini 2.5 Flash Lite. Beyond small systematic differences captured by the GLM, graders give very similar scores to individual items (pairwise Pearson correlations between graders is , ). Further, the grader model effect is substantially weaker than the effect of the base model class being steered. Smaller models generally seem to result in lower steering strengths than larger models. Figure˜14 displays the steering results for the different coherence graders. Across all graders, the key finding remains consistent: steering with the activation difference produces much higher similarity to the finetuning dataset than other sources.
D.3.1 Hypothesis Grader and Agent Interactions
Lastly, in Figure˜13 we show an extension to Figure˜6(c) that includes the effects of the interaction between the grader model and the agent model. We do not find robust evidence for interactions between hypothesis grader models and agent models: the posterior effects’ credible intervals all include zero.
The GLM study investigates whether there are systematic biases between graders. To assess inter-grader agreement, we compute Krippendorff’s [Krippendorff, 2018] with an ordinal distance metric across the three grader models. We obtain , indicating high agreement. To complement this, we compute Pearson correlations between grader pairs. All pairwise correlations are high (, ), indicating that the graders not only show strong overall agreement () but also rank items similarly.
Appendix E Generalization to broader finetuning
In this section, we demonstrate that the described phenomenon does not clearly manifest in more realistic, less narrowly-focused finetuning settings.
E.1 Chat Finetuning
We repeat our analysis by comparing the base and chat/instruct versions of Qwen3 1.7B, Llama3.2 1B, and Llama3.1 8B. Since we lack access to the true training dataset, we use tulu-3-sft-olmo-2-mixture [Lambert et al., 2025] as a proxy.
Figure˜15 shows the token relevance and steering results when analyzing chat finetuning across three models. We observe no clear difference between the metrics on the activation difference and the baseline, suggesting that the same type of trace either does not exist or is not as readily detectable. One notable exception is the token relevance for Qwen3 1.7B, which shows a clear difference between the activation difference and the baseline. Further investigation reveals that this stems from a single position where 7 Chinese tokens are tagged as relevant, displayed in Table˜2. While these tokens may be chat-relevant, their significance is difficult to determine without access to the training dataset to verify whether these phrases appear frequently. Table˜1 displays the relevant tokens for the of LLama3.2 1B, which also shows a slightly higher fraction of relevant tokens. However, these are primarily structural elements that we classify as false positives. We note that because we lack knowledge of the exact posttraining recipe used for these models, we provide a generic description of chat-tuning to the token relevance grader (see Prompt˜14).
| Token |
| package |
| import |
| def |
| Token | English Translation |
| 情况下 | under the circumstance |
| 状态下 | under the state/condition |
| 此基础上 | on this basis |
| 方式进行 | proceed in this manner |
| 基础上 | on the basis of |
| 环境中 | in the environment |
| 条件下 | under the condition |
E.2 More realistic domain finetuning and other modalities
In Figure˜17, we show the token relevance and steering results for the Domain organisms alongside the organisms we already analyzed. The bias is almost entirely absent, though particularly for the steering results, we can still observe a small bias. When comparing the Domain Finetuning self-sim results in Figure˜17(b) to the others Finetuning self-sim, we see that the dataset is much less narrow. Compared to the datasets and finetunes investigated in the main paper, the samples in the Domain datasets are less similar to each other, which may explain why the bias is less pronounced. In Figure˜16, we see that despite the small bias, the agent can still describe the finetuning objective in some cases. Notably, the ADL agent with clearly outperforms the strong baseline with interactions. Interestingly, the ADL agent with performs similarly as the baselines and mostly identifies that the finetuning involves visual instructions.
Appendix F Additional Results
F.1 Layer Ablation
We investigate how trace strength varies across different layers beyond the middle layer used in the main paper. We evaluate layers at , and depth (rounded down to the nearest layer) for Qwen3 1.7B and Llama3.2 1B Instruct on three SDF organisms (cake bake, fda approval, and kansas abortion). Figure˜18 shows the results, revealing two main patterns. First, token relevance results strengthen with layer depth, with the effect most pronounced in the last layer. This confirms that logit differences are also highly informative (see Aranguri and McGrath [2025]), which we leave for future work. Second, steering results deteriorate with layer depth, with the middle layer performing best. These findings confirm that the middle layer is optimal for our purposes: it provides the most effective steering while still yielding valuable information from the token results. We note that slightly deeper layers (e.g. depth) might perform even better, but we leave this exploration for future work.
F.2 Full Training Ablation
@@
Most investigated models are finetuned using LoRA. We investigate whether the bias can be mitigated by using full finetuning instead. We retrain Qwen3 1.7B, Llama3.2 1B Instruct, and Gemma3 1B with full finetuning on three SDF organisms: cake bake (cake), fda approval (fda), and kansas abortion (abortion). Figure˜19 shows the token and steering results for both full and LoRA finetuned models. Both training methods produce clearly detectable bias. Full finetuning shows slightly higher bias than LoRA finetuning. Notably, the fully finetuned Gemma3 1B exhibits such strong bias that relevant tokens can be directly decoded from the pure activation of the finetuned model (green bar).
F.3 Reducing Training Samples